find the characteristics of the equation pq=xy and determine the integral surface which passes through curve z=x, y=0
1
Expert's answer
2021-07-26T15:40:26-0400

Solving with method of characteristics :


"f(x, y, z, p, q)=pq-xy=0"

"f_x=-y, f_y=-x, f_z=0, f_p=q, f_q=p"

"\dfrac{dx}{f_p}=\dfrac{dy}{f_q}=\dfrac{dz}{pf_p+qf_q}=\dfrac{dp}{-(f_x+pf_z)}=\dfrac{dq}{-(f_y+qf_z)}"


"\dfrac{dx}{q}=\dfrac{dy}{p}=\dfrac{dz}{2pq}=\dfrac{dp}{y}=\dfrac{dq}{x}=dt"

Solve


"\dfrac{dp}{dt}=y, \dfrac{dy}{dt}=p"

"\dfrac{d^2p}{dt^2}=\dfrac{dy}{dt}"

"\dfrac{d^2p}{dt^2}=p"

"p=Ae^{t}+Be^{-t}"

"y=\dfrac{dp}{dt}=Ae^{t}-Be^{-t}"


"q=Ce^{t}+De^{-t}"

"x=\dfrac{dq}{dt}=Ce^{t}-De^{-t}"

"\dfrac{dz}{dt}=2pq=2(Ae^{t}+Be^{-t})(Ce^{t}+De^{-t})"

"=2ACe^{2t}+2BDe^{-2t}+2AD+2BC"

"z=ACe^{2t}-BDe^{-2t}+2ADt+2BCt+E"


"pq-xy=0=>(Ae^{t}+Be^{-t})(Ce^{t}+De^{-t})"

"-(Ae^{t}-Be^{-t})(Ce^{t}-De^{-t})=0"

"=>AD+BC=0"

The initial data curve is written in parametric form as


"x_0(s)=s, y_0(s)=0, z_0(s)=s"

Thus, using the initial data, the given PDE becomes


"p_0(s)q_0(s)=0"

The strip condition gives


"1=p_0(1)+q_0(0)"

Therefore we have


"x_0=s, y_0=0, z_0=s,p_0=1, q_0=0"

Substitute


"C-D=s"

"A-B=0"

"AC-BD+E=s"

"A+B=1"

"C+D=0"

"A=\dfrac{1}{2}, B=\dfrac{1}{2}, C=\dfrac{1}{2}s, D=-\dfrac{1}{2}s, E=\dfrac{1}{2}s"

Characteristics:


"x=\dfrac{e^t+e^{-t}}{2}s"

"y=\dfrac{e^t-e^{-t}}{2}"

"z=\dfrac{1}{2}(\dfrac{e^{2t}+e^{-2t}}{2})s+\dfrac{1}{2}"


"p=\dfrac{e^t+e^{-t}}{2}"

"q=\dfrac{e^t-e^{-t}}{2}s"





"x=s\text{cosh}t, y=\text{sinh}t"

"x^2=s^2\text{cosh}^2t , 1+y^2=1+\text{sinh}^2t=\text{cosh}^2t"

"z=\dfrac{1}{2}s(\text{cosh}(2t)+1)=s\text{cosh}^2t"

"z^2=s^2\text{cosh}^4t=x^2(1+y^2)"

Integral Surface: 


"z^2=x^2(1+y^2)"